Solve the problems below and show the complete solution.

What is the f of the pendulum on a clock with a L = 5.00 cm long?
f = (1 / 2π) √ (L / g)
1
Expert's answer
2020-12-14T04:39:54-0500

"f=\frac{1}{2\pi}\sqrt{g/L}=\frac{1}{2\cdot3.14}\sqrt{9.8/0.05}=2.23(Hz)"